# A long-lasting eruption heralds SN 2023ldh, a clone of SN 2009ip††thanks: Table A1 is only available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/XXX/AXXX

### v1

@@
A long-lasting eruption heralds SN 2023ldh, a clone of SN 2009ip††thanks: Table A1 is only available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/XXX/AXXX
We discuss the results of the spectroscopic and photometric monitoring of the type IIn supernova (SN) 2023ldh. Survey archive data show that the SN progenitor experienced some erratic outbursts in the years before exploding. From May 2023, the source shows a general slow luminosity rise lasting over four months with some superposed luminosity fluctuations. In analogy to SN~2009ip, we label this brightening as Event A. During Event A, SN~2023ldh reaches a maximum absolute magnitude of mag. Then the light curves show a luminosity decline of about 1 mag in all filters lasting about two weeks, followed by a steep brightening (Event B) to an absolute peak magnitude of mag, replicating the evolution of SN~2009ip and similar SNe IIn. Three spectra of SN~2023ldh are obtained during Event A, showing multi-component P Cygni profiles of H I and Fe II lines. During the rise to the Event B peak, the spectrum shows a blue continuum dominated by Balmer lines in emission with Lorentzian profiles, with a full width at half-maximum (FWHM) velocity of about 650 km s-1. Later, in the post-peak phase, the spectrum reddens, and broader wings appear in the H line profile. Metal lines are well visible with P Cygni profiles and velocities of about 2000 km s-1. Beginning around three months past maximum and until very late phases, the Ca II lines become among the most prominent features, while H is dominated by an intermediate-width component with a boxy profile. Although SN~2023ldh mimics the evolution of other SN~2009ip-like transients, it is slightly more luminous and has a slower photometric evolution. The surprisingly homogeneous observational properties of SN~2009ip-like events may indicate similar explosion scenarios and similar progenitor parameters.
@@
Spectroscopic monitoring of the pre-SN eruptive phase is available in a handful of SNe IIn: SN~2009ip (Pastorello et al., 2013), SN~2015bh (Elias-Rosa et al., 2016; Thöne et al., 2017), SN~2016jbu (Kilpatrick et al., 2017; Brennan et al., 2022a). More recently, pre-SN spectra were also obtained for the type Ibn SN~2023fyq (Brennan et al., 2024; Dong et al., 2024). The combination of photometric monitoring and spectroscopic observations obtained before the stellar core collapse is fundamental for the characterization of the progenitor star. For this reason, it is crucial to obtain spectral sequences for all species of transients sitting in the luminosity gap that separates classical novae and core-collapse supernovae (Kasliwal, 2012; Pastorello & Fraser, 2019; Cai et al., 2022), as a fraction of them have been seen to explode as a SN a short time later. In this paper, we discuss the case of an object initially identified as a long-lasting gap transient that later experienced a major brightening compatible with a type IIn SN explosion: SN~ 2023ldh. A combination of long-lasting photometric monitoring and spectra from the pre-SN eruptive phase to the late nebular phase makes SN~2023ldh one of the interacting SNe with the most comprehensive datasets.
The optical transient SN~2023ldh111The object is also known with the following survey designations: ZTF23aamanim, ATLAS23qiw and PS23hmg. was discovered by the Zwicky Transient Facility (ZTF, Bellm et al., 2019; Graham et al., 2019) on 2023 May 28.32 UT, at a -band magnitude of 20.74 (De, 2023). The coordinates are: RA = 15:09:09.597 and Dec = +52:31:59.80 (J2000), west and north of the centre of the spiral galaxy (SAb-type) NGC~5875. The location of SN~2023ldh is nearly coincident with an extended source, SDSS J150909.51+523159.0, and is also very close (about 0.2 arcmin) to the explosion site of a recent, peculiar type Ic event, SN~2022oqm (Sollerman et al., 2022; Zimmerman et al., 2022; Fulton et al., 2022). The closest source within NGC~5875 with known redshift (z ) is SDSS J150912.08+523154.3, only 0.39 arcmin south-east of SN~2023ldh.
@@ 2 Observations
A short time after the discovery announcement of SN~2023ldh, we initiated a monitoring campaign aimed to study the evolution of the transient in the context of the Nordic Optical Telescope (NOT) Unbiased Transient Survey-2 (NUTS-2) program.333https://nuts.sn.ie/ For the photometric campaign, we made use of the following facilities: the 2m Liverpool Telescope (LT) equipped with IO:O, the 2.56m Nordic Optical Telescope (NOT) with the Alhambra Faint Object Spectrograph and Camera (ALFOSC), the 10.4m Gran Telescopio Canarias (GTC) with the Optical System for Imaging and low-Intermediate-Resolution Integrated Spectroscopy (OSIRIS+), hosted in La Palma (Spain); the 1.82m Copernico Telescope with the Asiago Faint Object Spectrograph and Camera (AFOSC) and the 67/91cm Schmidt Telescope with a Moravian detector, hosted at the Asiago Observatory, Mt. Ekar (Italy); the Wide-field Optical Telescope (WOT), a 68/91cm Schmidt Telescope equipped with an Apogee Aspen CCD, at the Campo Imperatore Observatory (Italy); the Joan Oró Telescope (TJO) with MEIA2 at the Montsec Astronomical Observatory (Spain); the 1m-class telescopes of Las Cumbres Observatory (LCO) telescope network444https://lco.global/ equipped with Sinistro cameras in the framework of the Global Supernova Project. Additional data were collected through 0.5m class amateur facilities of the iTelescope network555https://support.itelescope.net/support/home hosted in Utah and California (USA), and others were obtained at Itagaki Observatory of Yamagata (Japan). Crucial pre-SN data were finally collected from the public surveys, including ZTF 666Images retrieved through the NASA/IPAC Infrared Science Archive: https://irsa.ipac.caltech.edu/Missions/ztf.html. (in the and bands), ATLAS777The template-subtracted forced photometry is released through the Asteroid Terrestrial-impact Last Alert System (ATLAS) data-release interface: https://fallingstar-data.com/forcedphot/queue/. (with the cyan and orange filters – hereafter and ; Tonry et al., 2018; Smith et al., 2020), and Pan-STARRS (in the , , and bands; Chambers et al., 2019; Flewelling et al., 2020; Magnier et al., 2020). Including the survey data, the photometric data analysed in this paper cover a period of about nine years, from 2016 to early 2025.
@@ 3.1 Light curve Evolution
Earlier observations, obtained from mid-2016 by the main surveys or through the inspection of the public archives, show no sources to a limiting magnitude of about 22 in the band (see Table A.1). However, Pan-STARRS observations of the SN field show sparse detections in the band in May 2018, July-August 2020 and April 2021 (with -band magnitudes ranging from 19 to 22 mag). This suggests a former, erratic phase of variability for the same stellar system that later generated SN 2023ldh, similar to the extreme variability shown by massive binaries containing luminous blue variables (LBVs; e.g., Pastorello et al., 2010; Müller et al., 2023; Aghakhanloo et al., 2023).
The transient was then observed again starting from 2023 May 9, about 20 days before the official discovery date (De, 2023). In this phase, the light curve of the transient shows a slow increase in luminosity (by about 2.5-3 magnitudes) in all optical filters lasting about four months. Significant luminosity fluctuations are observed superimposed on the global luminosity rise, in particular at about 1.5 months after the discovery. The average rise rates are mag (100 d)-1 in the band, mag (100 d)-1 in the band, and mag (100 d)-1 in the band. The epoch and the magnitude of the -band light-curve maximum (MJD and mag, respectively) are determined through a low-order polynomial fit. This first peak is reached about four months after the discovery, at a similar time for all filters (see Table 1). Accounting for the line-of-sight extinction and the adopted host galaxy distance, we obtain the following absolute magnitude at the time of this first light-curve peak: mag.111111The error on the absolute magnitude is largely dominated by the uncertainty on the host galaxy distance. Following the notation introduced by Pastorello et al. (2013), we label this early peak as Event A. We note a moderately blue colour at the time of the Event A maximum, mag.
@@ 4.2 Spectral Evolution
Rise to Event B maximum – After the Event B onset, on MJD , the continuum temperature rises and reaches 16000-17000 K a few days before the light-curve maximum, then gradually declines in the post-maximum evolution. We note that the spectral energy distribution (SED) peaks in the ultra-violet (UV) domain at these early epochs. Since this region is not covered by our optical spectra, the temperature estimates at the early phases obtained via blackbody fits to the spectral continuum are affected by large errors (typically 2000-3000 K). During the rise to the Event B peak (which lasts 15 days) and up to about three weeks past maximum, the SN spectrum shows very little evolution. It is characterized by a blue continuum with superposed prominent Balmer lines showing a dominant emission component. The line profile is best fitted with a Lorentzian function whose emission core is at the rest wavelength of the transition. The FWHM velocity of the H lines is km s-1, as inferred from the best-resolution ( km s-1) GTC/OSIRIS+ spectrum obtained d from the maximum light. A residual broad P Cygni absorption is detected at the positions of H and H with a minimum blueshift of about 4000 km s-1 and a blue wing that extends to about 8000-10000 km s-1. We also detect He I lines (4471.5, 4921.9, 5015.7, 5875.6, 6678.1, 7065.2), revealing a narrow emission component superposed on a broad P Cygni base. While the FWHM velocity of the emission component is comparable to that of the narrow Balmer lines, He I 5875.6 shows a clear P Cygni profile, with a minimum at km s-1 from the rest frame, and a blue-shifted wing extending up to about 5000 km s-1. Another emission feature clearly detected in the low-resolution spectra from d to d from maximum is O I 8446.5, with 400-450 km s-1. Spectra showing a prominent O I 8446.5 emission line, while O I 7772-7775 remains below the detection threshold, are quite frequent among SNe IIn (e.g., in SN~1995G; Pastorello et al., 2002). This feature is likely indicative of Ly pumping (see Valerin et al., 2025, for a detailed discussion on the Bowen fluorescence mechanism). Finally, a broad emission bump is observed at 4600-4700 Å, possibly due to a blend of C III, N III and He II lines.
@@ 6 Comparison with other SN 2009ip-like events
In Fig. 5 we compare the observed properties of SN~2023ldh with those of photometrically similar SNe IIn for which Event A was well sampled both in photometry and spectroscopy: SN~2009ip, SN~2015bh (Elias-Rosa et al., 2016; Goranskij et al., 2016; Thöne et al., 2017) and SN~2016jbu (Brennan et al., 2022a, b). The type Ibn SN~2023fyq (Brennan et al., 2024; Dong et al., 2024) is also considered as a comparison object, although spectroscopically different and with a lower-luminosity Event A than other objects of the sample. The “optical” pseudo-bolometric light curves of the five objects are shown in the top-left panel, while the other three panels show their optical spectra at three reference phases: during the rise to the Event A maximum (phase d; top right), in the proximity of the Event B peak (phases to d; bottom-left), and at late phases ( d; bottom-right).
@@
Other ejecta-CSM interacting SNe share similarities with the above objects, including several type IIn SNe191919The group of SN 2009ip-like type IIn SNe includes also SN~2010mc (Ofek et al., 2013), SN 2013gc (Reguitti et al., 2019), SNhunt151 (Elias-Rosa et al., 2018), SN 2016cvk (Matilainen et al., 2025), SN~2018bdu (Pastorello et al., 2018), SN~2018cnf (Pastorello et al., 2019), SN 2019zrk (Fransson et al., 2022), SN 2021qqp (Hiramatsu et al., 2024) and the 2024 event following SN 2022mop (Brennan et al., 2025). and two transitional type IIn/Ibn events, SN~2021foa (Reguitti et al., 2022; Farias et al., 2024; Gangopadhyay et al., 2025) and SN~2022pda (Cai et al., 2025). However, in all these cases, the available spectroscopic and photometric data do not completely cover the crucial phases of the SN evolution illustrated in Fig. 5.
@@ 7 Eruption or Supernova Explosion?
Long-lasting eruptive phases accompanied by a global trend of slow luminosity rise are typical of the so-called luminous red novae (LRNe, e.g., Pastorello et al., 2023, and references therein), that are thought to result from merging events of non-degenerate stars. Schrøder et al. (2020) suggested that some SNe IIn are the outcome of the coalescence of a compact object, a black hole (BH) or a neutron star (NS), with the core of a massive companion, while the system is embedded in a H-rich common envelope.
@@
Recently, Brennan et al. (2025) presented a detailed study on SN~2022mop, a transient source characterized by two spatially coincident outbursts, both with a duration and a luminosity consistent with those of a SN explosion. The first outburst (in mid-2022) was poorly observed and spectroscopically classified a few months later as a stripped-envelope SN in the nebular phase. The second outburst (in late 2024), better sampled, had instead the spectrum of a young type IIn and a double-peak light curve similar to SN~2009ip and SN~2023ldh. Between the two main events, a few faint ( mag) short-duration outbursts were also observed (e.g., in October 2023). Brennan et al. (2025) proposed the 2022 event to be a genuine stripped-envelope SN in a binary system, whose remnant - likely a NS in an unstable eccentric orbit - started to interact with the outer layers of the H-rich secondary producing the recurrent flares observed in 2023. The interaction within a common envelope led the orbital period to gradually decrease, and the NS to merge into the secondary producing the SN~2009ip-like event observed in 2024.
@@
A somewhat similar scenario is proposed by Tsuna et al. (2024) for the type Ibn SN~2023fyq, in which the observed properties of the SN are explained invoking a binary system formed by a He donor star transferring mass to a NS (or a BH) companion. The super-Eddington mass transfer triggers a circum-binary outflow from the Lagrangian point L2 and a fast wind from the accretion disk of the degenerate companion, which power the slow-rising light curve of SN~2023fyq in the years preceding the main event. The consequent loss of angular momentum leads the two components to merge, producing an outburst powered by ejecta-CSM interaction with the properties of a type Ibn SN, without necessarily invoking a core collapse. According to Tsuna et al. (2024), the above model may only explain transients showing a smooth pre-outburst light-curve rise. For example, while the scenario is applicable to the evolution of the type IIn SN~2021qqp (Hiramatsu et al., 2024) and objects such as SNhunt151 (Elias-Rosa et al., 2018), it cannot comfortably explain the interacting SNe heralded by an erratic and fast-evolving variability (but, see Brennan et al., 2025).
@@
The rapid photometric variability followed by the slow photometric rise typical of SN~2009ip-like events is somewhat reminiscent of the pre-outburst phase in LRNe (e.g., Soker & Kashi, 2013). The long-term OGLE light curve of LRN V1309 Sco (Tylenda et al., 2011) is characterized by multiple distinct phases: i. a slow rise lasting several years with a superposed modulations characteristic of a binary system; ii. an initial luminosity decline followed by a rebrightening lasting a few months is expected when the binary system is engulfed in an optically thick common envelope, but it is still losing mass; iii. the double-peaked outburst due to the merging event and the subsequent gas outflow. Whether the rapid luminosity oscillations observed before Events A in SN~2009ip-like events, including those observed in SN~2023ldh from years to about year before the outbursts, are indicative of a modulated light curve due to binary interaction, or the erratic variability similar to that expected in an LBV202020An example is V37 in NGC 2403 (Weis & Bomans, 2005; Maund et al., 2006). is still unclear. Well-cadenced data with high S/N are required to eventually reveal periodicity in the archival observation. However, we note that some periodicity in the mass-loss events suggesting binary interaction was revealed by the bumps in the light curve of Event B in SN~2009ip (Kashi et al, 2013; Martin et al., 2015).
@@
Eruptive wave-driven outbursts (e.g., Quataert & Shiode, 2012; Moriya et al., 2014; Leung & Fuller, 2020; Chugai, 2022) or steady wind mass-loss resulting from a continuous energy deposition at the base of the stellar envelope in progenitors exceeding the Eddington’s luminosity are still plausible explanations for the pre-SN evolution of SN~2023ldh and similar events (Matsumoto et al., 2022). In this case, the mass loss would generate an optically thick CSM engulfing the progenitor. The subsequent stellar core-collapse ejects high-velocity material which interacts with the CSM, sustaining the light curve during Event B. We remark, however, that this scenario cannot comfortably reproduce the frequent short-duration outbursts below absolute magnitude of observed in the archival light curves of many SN~2009ip-like transients.
@@
Another scenario was proposed for SN~2009ip-like objects, according to which the entire light curve is explained through a non-terminal eruptive phase of a very massive star. In particular, the CSM is produced after a long-lasting (years) period of erratic variability, during which the star loses a significant amount of envelope mass. Later, the star experiences a giant eruption (Event A) and the interaction between the material ejected in Event A with the previously gathered CSM powers Event B. So, in this case, the terminal SN explosion is not required for reproducing the SN observables. Of course, only very massive stars can release enough mass and radiate such high energy for a long time without the need to invoke a terminal SN explosion. This is, for instance, proposed in the pulsational pair-instability scenario (Woosley, 2017).

### v2

@@
A long-lasting eruption heralds SN 2023ldh, a clone of SN 2009ip
We discuss the results of the spectroscopic and photometric monitoring of the type IIn supernova (SN) 2023ldh. Survey archive data show that the SN progenitor experienced erratic variability in the years before exploding. From May 2023, the source shows a general slow luminosity rise lasting over four months with some superposed luminosity fluctuations. In analogy to SN~2009ip, we label this brightening as Event A. During Event A, SN~2023ldh reaches a maximum absolute magnitude of mag. Then the light curves show a luminosity decline of about 1 mag in all filters lasting about two weeks, followed by a steep brightening (Event B) to an absolute peak magnitude of mag, replicating the evolution of SN~2009ip and similar SNe IIn. Three spectra of SN~2023ldh are obtained during Event A, showing multi-component P Cygni profiles of H I and Fe II lines. During the rise to the Event B peak, the spectrum shows a blue continuum dominated by Balmer lines in emission with Lorentzian profiles, with a full width at half-maximum (FWHM) velocity of about 650 km s-1. Later, in the post-peak phase, the spectrum reddens, and broader wings appear in the H line profile. Metal lines are well visible with P Cygni profiles and velocities of about 2000 km s-1. Beginning around three months past maximum and until very late phases, the Ca II lines become among the most prominent features, while H is dominated by an intermediate-width component with a boxy profile. Although SN~2023ldh mimics the evolution of other SN~2009ip-like transients, it is slightly more luminous and has a slower photometric evolution. The surprisingly homogeneous observational properties of SN~2009ip-like events may indicate similar explosion scenarios and similar progenitor parameters.
@@
Spectroscopic monitoring of the pre-SN eruptive phase is available in a handful of SNe IIn: SN~2009ip (Smith et al., 2010; Foley et al., 2011; Pastorello et al., 2013; Mauerhan et al., 2013), SN~2015bh (Elias-Rosa et al., 2016; Thöne et al., 2017), SN~2016jbu (Kilpatrick et al., 2017; Brennan et al., 2022a). More recently, pre-SN spectra were also obtained for the type Ibn SN~2023fyq (Brennan et al., 2024; Dong et al., 2024). The combination of photometric monitoring and spectroscopic observations obtained before the stellar core collapse is fundamental for the characterization of the progenitor star. For this reason, it is crucial to obtain spectral sequences for all species of transients sitting in the luminosity gap that separates classical novae and core-collapse supernovae (Kasliwal, 2012; Pastorello & Fraser, 2019; Cai et al., 2022), as a fraction of them have been seen to explode as a SN a short time later. In this paper, we discuss the case of an object initially identified as a long-lasting gap transient that later experienced a major brightening compatible with a type IIn SN explosion: SN~ 2023ldh. A combination of long-lasting photometric monitoring and spectra from the pre-SN eruptive phase to the late nebular phase makes SN~2023ldh one of the interacting SNe with the most comprehensive datasets.
The optical transient SN~2023ldh111The object is also known with the following survey designations: ZTF23aamanim, ATLAS23qiw and PS23hmg. was discovered by the Zwicky Transient Facility (ZTF, Bellm et al., 2019; Graham et al., 2019) on 2023 May 28.32 UT, at a -band magnitude of 20.74 (De, 2023). The coordinates are: RA = 15:09:09.597 and Dec = +52:31:59.80 (J2000), west and north of the centre of the spiral galaxy (SAb-type) NGC~5875. The location of SN~2023ldh is nearly coincident with an extended source, SDSS J150909.51+523159.0, and is also very close (about 0.2 arcmin) to the explosion site of the Ca-rich transient SN~2022oqm (Sollerman et al., 2022; Zimmerman et al., 2022; Fulton et al., 2022; Irani et al., 2024; Yadavalli et al., 2024). The closest source within NGC~5875 with known redshift (z ) is SDSS J150912.08+523154.3, only 0.39 arcmin south-east of SN~2023ldh.
@@ 2 Observations
A short time after the discovery announcement of SN~2023ldh, we initiated a monitoring campaign aimed to study the evolution of the transient in the context of the Nordic Optical Telescope (NOT) Unbiased Transient Survey-2 (NUTS-2) program.333https://nuts.sn.ie/ For the photometric campaign, we made use of the following facilities: the 2m Liverpool Telescope (LT) equipped with IO:O, the 2.56m Nordic Optical Telescope (NOT) with the Alhambra Faint Object Spectrograph and Camera (ALFOSC), the 10.4m Gran Telescopio Canarias (GTC) with the Optical System for Imaging and low-Intermediate-Resolution Integrated Spectroscopy (OSIRIS+), hosted in La Palma (Spain); the 1.82m Copernico Telescope with the Asiago Faint Object Spectrograph and Camera (AFOSC) and the 67/91cm Schmidt Telescope with a Moravian detector, hosted at the Asiago Observatory, Mt. Ekar (Italy); the Wide-field Optical Telescope (WOT), a 68/91cm Schmidt Telescope equipped with an Apogee Aspen CCD, at the Campo Imperatore Observatory (Italy); the Joan Oró Telescope (TJO) with MEIA2 at the Montsec Astronomical Observatory (Spain); the 1m-class telescopes of Las Cumbres Observatory (LCO) telescope network444https://lco.global/ equipped with Sinistro cameras in the framework of the Global Supernova Project. Additional data were collected through 0.5m class amateur facilities of the iTelescope network555https://support.itelescope.net/support/home hosted in Utah and California (USA), and others were obtained at Itagaki Observatory of Yamagata (Japan). Crucial pre-SN data were finally collected from the public surveys, including ZTF 666Images retrieved through the NASA/IPAC Infrared Science Archive: https://irsa.ipac.caltech.edu/Missions/ztf.html. (in the and bands), ATLAS777The template-subtracted forced photometry is released through the Asteroid Terrestrial-impact Last Alert System (ATLAS) data-release interface: https://fallingstar-data.com/forcedphot/queue/. (with the cyan and orange filters – hereafter and ; Tonry et al., 2018; Smith et al., 2020), and Pan-STARRS (in the , , and bands; Chambers et al., 2016; Flewelling et al., 2020; Magnier et al., 2020). Including the survey data, the photometric data analysed in this paper cover a period of about nine years, from 2016 to early 2025.
@@ 3.1 Light curve evolution
Earlier observations, obtained from mid-2016 by the main surveys or through the inspection of the public archives, show no sources to a limiting magnitude of about 22 in the band (see Table A.1). Subsequent observations return in most cases detection limits, although a faint source is detected () in the SN location at several epochs. In particular, the source is visible in -band Pan-STARRS images obtained in mid-2019 (from May 20 to July 18), at magnitudes ranging from to mag. In -band Pan-STARRS and ATLAS images taken from April to September 2020 and from January to July 2021, the object is seen always at magnitudes fainter than mag. In January and February 2022 the source is still detected in ATLAS images at slightly brighter magnitudes ( mag). The source is also found in archival LCOGT images obtained in mid-2022 (on July 18, August 27, and September 13) with magnitudes varying erratically in the three filters (from to mag in the band). These luminosity fluctuations are produced by the same stellar system that will later generate SN 2023ldh, and are similar to the extreme variability expected in massive binaries containing luminous blue variables (LBVs; e.g., Wagner et al., 2004; Pastorello et al., 2010; Müller et al., 2023; Aghakhanloo et al., 2023a, b; Pustilnik & Perepelitsyna, 2025).
The transient is observed again from 2023 May 9, about 20 days before the official discovery date (De, 2023). In this phase, the light curve of the transient shows a slow increase in luminosity (by about 2.5-3 magnitudes) in all optical filters lasting about four months. Significant luminosity fluctuations are observed superimposed on the global luminosity rise, in particular at about 1.5 months after the discovery. The average rise rates are mag (100 d)-1 in the band, mag (100 d)-1 in the band, and mag (100 d)-1 in the band. The epoch and the magnitude of the -band light-curve maximum (MJD and mag, respectively) are determined through a low-order polynomial fit. This first peak is reached about four months after the discovery, at a similar time for all filters (see Table 1). Accounting for the line-of-sight extinction and the adopted host galaxy distance, we obtain the following absolute magnitude at the time of this first light-curve peak: mag.111111The error on the absolute magnitude is largely dominated by the uncertainty on the host galaxy distance. Following the notation introduced by Pastorello et al. (2013), we label this early peak as Event A. We note a moderately blue colour at the time of the Event A maximum, mag.
@@ 4.2 Spectral evolution
Rise to Event B maximum – After the Event B onset, on MJD , the continuum temperature rises and reaches 16000-17000 K a few days before the light-curve maximum, then gradually declines in the post-maximum evolution. We note that the spectral energy distribution (SED) peaks in the ultra-violet (UV) domain at these early epochs. Since this region is not covered by our optical spectra, the temperature estimates at the early phases obtained via blackbody fits to the spectral continuum are affected by large errors (typically 2000-3000 K). During the rise to the Event B peak (which lasts 15 days) and up to about three weeks past maximum, the SN spectrum shows very little evolution. It is characterized by a blue continuum with superposed prominent Balmer lines showing a dominant emission component. The line profile is best fitted with a Lorentzian function whose emission core is at the rest wavelength of the transition. The FWHM velocity of the H lines is km s-1, as inferred from the best-resolution ( km s-1) GTC/OSIRIS+ spectrum obtained d from the maximum light. A residual broad P Cygni absorption is detected at the positions of H and H with a minimum blueshift of about 4000 km s-1 and a blue wing that extends to about 8000-10000 km s-1. We also detect He I lines (4471.5, 4921.9, 5015.7, 5875.6, 6678.1, 7065.2), revealing a narrow emission component superposed on a broad P Cygni base. While the FWHM velocity of the emission component is comparable to that of the narrow Balmer lines, He I 5875.6 shows a clear P Cygni profile, with a minimum at km s-1 from the rest frame, and a blue-shifted wing extending up to about 5000 km s-1. Another emission feature clearly detected in the low-resolution spectra from d to d from maximum is O I 8446.5, with 400-450 km s-1. Spectra showing a prominent O I 8446.5 emission line, while O I 7772-7775 remains below the detection threshold, are quite frequent in SNe IIn, such as SN~1995G (Pastorello et al., 2002) and SN~2010jl (Fransson et al., 2014). This feature is likely indicative of Ly pumping (see Valerin et al., 2025, for a detailed discussion on the Bowen fluorescence mechanism). Finally, a broad emission bump is observed at 4600-4700 Å, possibly due to a blend of C III, N III and He II lines.
@@ 6 Comparison with other SN 2009ip-like events
In Fig. 5 we compare the observed properties of SN~2023ldh with those of photometrically similar SNe IIn for which Event A was well sampled both in photometry and spectroscopy: SN~2009ip, SN~2015bh (Elias-Rosa et al., 2016; Goranskij et al., 2016; Ofek et al., 2016; Thöne et al., 2017) and SN~2016jbu (Brennan et al., 2022a, b). The type Ibn SN~2023fyq (Brennan et al., 2024; Dong et al., 2024) is also considered as a comparison object, although spectroscopically different and with a lower-luminosity Event A than other objects of the sample. The “optical” pseudo-bolometric light curves of the five objects are shown in the top-left panel, while the other three panels show their optical spectra at three reference phases: during the rise to the Event A maximum (phase d; top right), in the proximity of the Event B peak (phases to d; bottom-left), and at late phases ( d; bottom-right).
@@
Other ejecta-CSM interacting SNe share similarities with the above objects, including several type IIn SNe191919The group of SN 2009ip-like type IIn SNe includes also SN~2010mc (Ofek et al., 2013), iPTF13z (Nyholm et al., 2017), SN 2013gc (Reguitti et al., 2019), SNhunt151 (Elias-Rosa et al., 2018), LSQ13zm (Tartaglia et al., 2016), SN 2016cvk (Matilainen et al., 2025), SN~2016bdu (Pastorello et al., 2018), SN~2018cnf (Pastorello et al., 2019), SN 2019zrk (Fransson et al., 2022), SN 2021qqp (Hiramatsu et al., 2024) and the 2024 event following SN 2022mop (Brennan et al., 2025). and two transitional type IIn/Ibn events, SN~2021foa (Reguitti et al., 2022; Farias et al., 2024; Gangopadhyay et al., 2025) and SN~2022pda (Cai et al., in prep.). However, in all these cases, the available spectroscopic and photometric data do not completely cover the crucial phases of the SN evolution illustrated in Fig. 5.
@@ 7 Eruption or supernova explosion?
Long-lasting eruptive phases accompanied by a global trend of slow luminosity rise are typical of the so-called luminous red novae (LRNe, e.g., Pastorello et al., 2023, and references therein), which are thought to result from merging events of non-degenerate stars. For example, SN~2009ip-like events show rapid photometric variability followed by a slow brightening, somewhat reminiscent of the pre-outburst phase of LRNe (e.g., V1309 Sco; Tylenda et al., 2011; Soker & Kashi, 2013).
@@
Schrøder et al. (2020) suggested that some SNe IIn are the outcome of the coalescence of a compact object, a black hole (BH) or a neutron star (NS), with the core of a massive companion, while the system is embedded in a H-rich common envelope. Brennan et al. (2025) presented a study on SN~2022mop, a transient source characterized by two spatially coincident outbursts, both with duration and luminosity consistent with a SN explosion. Brennan et al. (2025) proposed that the first event in 2022 was a genuine stripped-envelope SN in a binary system, whose remnant - likely a NS in an unstable eccentric orbit - started to interact with the outer layers of the H-rich secondary producing the photometric fluctuations observed in 2023. The interaction within a common envelope led the orbital period to gradually decrease, and the NS to merge into the secondary producing a SN~2009ip-like event in 2024.
@@
A somewhat similar scenario is proposed by Tsuna et al. (2024) for the type Ibn SN~2023fyq, whose properties favour a binary system with a He donor star transferring mass to a NS (or BH) companion. The super-Eddington mass transfer triggers a circum-binary outflow from the Lagrangian point L2 and a fast wind from the accretion disk of the degenerate companion, which power the slow-rising light curve of SN~2023fyq in the years preceding the main event. The loss of angular momentum leads the two stars to merge, producing an outburst with the properties of an SN Ibn, without necessarily a terminal core collapse. According to Tsuna et al. (2024), the above model may only explain transients with smooth brightenings. Hence, while the scenario is applicable to the evolution of SNe IIn such as SN~2021qqp (Hiramatsu et al., 2024) and SNhunt151 (Elias-Rosa et al., 2018), it cannot comfortably explain SN~2009ip-like events characterized by former erratic, fast-evolving variability (but, see Brennan et al., 2025).
@@
Whether the rapid luminosity fluctuations of SN~2023ldh from to years before Event A are indicative of binary interaction or erratic variability similar to LBVs202020An example is V37 in NGC 2403 (Weis & Bomans, 2005; Maund et al., 2006; Humphreys et al., 2017). is unclear. Well-cadenced data with high S/N are required to eventually reveal periodicity in the archival observation. However, we note that some periodicity in the mass-loss events suggesting binary interaction was revealed by the bumps in the light curve of Event B in SN~2009ip (Kashi et al, 2013; Martin et al., 2015).
@@
Eruptive wave-driven outbursts (e.g., Quataert & Shiode, 2012; Moriya et al., 2014; Leung & Fuller, 2020; Chugai, 2022) or steady wind mass-loss resulting from a continuous energy deposition at the base of the stellar envelope in progenitors exceeding the Eddington’s luminosity are still plausible explanations for the pre-SN evolution of SN~2023ldh and similar events (Matsumoto et al., 2022). Mass loss generates an optically thick CSM engulfing the progenitor. The subsequent stellar core-collapse triggers high-velocity ejecta which collide with the CSM, sustaining the Event B light curve. However, this scenario cannot reproduce the frequent short-duration flares at observed before Event A.
@@
Another scenario was proposed for SN~2009ip-like objects, according to which the entire light curve is explained through a non-terminal eruptive phase of a very massive star. In particular, the CSM is produced after a long-lasting (years) period of erratic variability, during which the star loses a significant amount of envelope mass. Later, the star experiences a giant eruption (Event A) and the interaction between the material ejected in Event A with the previously gathered CSM powers Event B. In this case, the terminal SN explosion is not required for reproducing the SN observables. However, only very massive stars can release enough mass and radiate such high energy for a long time without the need to invoke a terminal explosion (e.g., in a pulsational pair-instability event; Woosley, 2017).
@@
Data availability
Table A.1 only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.